OpenStudy (anonymous):

How many solutions exist for the given equation? 12x + 1 = 3(4x + 1) – 2

OpenStudy (anonymous):

Do you know where to start? @ohgee

OpenStudy (anonymous):

12x+1=12x+3-2 12x-12x=1-1 0 Endless solution, because the equation doesn't depend on the x value .

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

Please dont put that its not right :( sorry

OpenStudy (anonymous):

@ohgee

OpenStudy (anonymous):

okay okay .. is it zero?

OpenStudy (anonymous):

nvm its right sry :P

OpenStudy (anonymous):

ok